# Echoing: Identity Failures when LLM Agents Talk to Each Other

### v1

@@ Abstract
As large language model (LLM) based agents interact autonomously with one another, a new class of failures emerges that cannot be predicted from single agent performance: behavioral drifts in agent-agent conversations (AxA). Unlike human–agent interactions, where humans ground and steer conversations, AxA lacks such stabilizing signals, making these failures unique. We investigate one such failure, echoing, where agents abandon their assigned roles and instead mirror their conversational partners, undermining their intended objectives. Through experiments across $60$ AxA configurations, $3$ domains, and $2000+$ conversations, we demonstrate that echoing occurs across three major LLM providers, with echoing rates from $5\%$ to $70\%$ depending on the model and domain. Moreover, we find that echoing is persistent even in advanced reasoning models with substantial rates ($32.8\%$) that are not reduced by increased reasoning efforts. We analyze prompt impacts, conversation dynamics, showing that echoing arises as interaction grows longer ($7+$ turns in experiments) and is not merely an artifact of sub-optimal prompting. Finally, we introduce a protocol-level mitigation in which targeted use of structured responses reduces echoing to $9\%$.
@@ 1 Introduction
Recent advances in large language models (LLMs) jaech2024openai; comanici2025gemini; guo2025deepseek have enabled agentic systems that can plan, reason, and act in open-ended settings (mialon2023augmented; guo2024large; wang2024survey). A natural next step is agent-agent (AxA) interaction: agents that converse directly with one another to collaborate, negotiate, and execute tasks on behalf of users or organizations (raskar2025beyond; beeai2025acp; google2025a2a; cisco2025Agntcy; tomasev2025virtual). Although AxA presents a promising frontier, our understanding of its reliability and requirements is still incomplete; including how consistently agents maintain their roles and objectives across multi-turn interactions.
@@
Existing research and evaluation frameworks are focused on single-agent capabilities, i.e., how well a model performs tasks in isolation or in a human supervised settings (yao2024tau; huang2025crmarena; sirdeshmukh2025multichallenge). Even when multi-turn dialog is considered, benchmarks typically rely on simulated user (instructions, tools, or context) and measure single-agent success criteria (barres2025tau). These task driven setups are unable to capture behaviors that emerge specifically from AxA, where agents with private, potentially misaligned, objectives must interact, namely, an user agent that is not simulated only to provide information for task completion. In contrast to human–agent interactions, where human feedback is accumulated continually, and often subtly, steering and grounding behavior, AxA relies on predefined instructions/specifications and an often untested assumption of alignment between agents (see Figure 1).
@@
This paper studies an AxA specific failure mode that we call echoing: an agent abandons its assigned identity and mirrors its conversational partner, undermining its stated objectives. We provide a systematic investigation of this behavioral failure across several models, domains, and prompts. Through experiments spanning $60$ AxA configurations, $3$ domains (car sales, hotel booking, supply chain), and $2000+$ conversations, we find that echoing occurs across three major providers with rates ranging from $5\%$ to $70\%$ depending on the model and domain. Echoing persists even in advanced reasoning models (average $32.8\%$) and does not diminish with increased reasoning effort; non-reasoning variants average $37.7\%$ (section 4.2.1). Our empirical results shows that prompt engineering reduces but does not eliminate the failure (section 4.2.2), suggesting a fundamental limitation akin to hallucination but specific to AxA. We further observe that standard completion metrics mask these failures as $93\%$ of conversations completed successfully, even when identity drift occurred. Further, the outcome value varied substantially within a single setup.
@@
Figure 1: Agent x Agent setup: (Left) Two agents, a customer agent and a seller agent, given instructions, objectives, private tools, and resources, are entrusted to complete a particular task given a situation-specific spec. The customer agent, in this case, negotiates a room on behalf of a human (with specific requirements) with a seller agent, a hotel agent representing an enterprise with specifics pertinent to the hotel. (Right) Conversation snippet from an AxA exchange where the customer agent echoes the language and behavior more appropriate of an hotel agent. The seller agent, in this example, continued the interaction without correction and ended up accepting the package proposed by the customer agent. Such a failure is unlikely in human–agent interactions and even when it arises would typically be corrected by the human ensuring that the agent remains aligned with its intended role. More examples of echoing are provided in Appendix D.
@@
Our setting differs from multi-agent systems (MAS) (stone2000multiagent; tampuu2017multiagent) in that AxA agents maintain private internal state, operate with distinct tools, and may have competing utilities. In contrast, MAS emphasize coordinated teams with shared goals, centralized orchestration, and even shared states; recent LLM-based MAS continues this paradigm for task decomposition and parallelism (wu2023autogen; hong2023metagpt; chen2024agentverse). This distinction is crucial as the real promise of AxA is that agents will act on our behalf out in the world, not orchestrate to complete a single turn of interaction with a human. In conversational AI, existing works study human–agent interaction where humans provide feedback and course correction (serban2017deep; roller2020recipes). Moreover, post-training, alignment techniques are optimized for human-facing use (ouyang2022training; anthropic2022constitutional; rafailov2023direct; song2024preference) and not AxA (might even be cause of biases such as over-accommodation, role drifts). Work on alignment, both objective adherence (christiano2017deep; russell2019human) and multi-agent coordination (carroll2019utility; eccles2019biases), rarely addresses an AxA setting. Other works on emergent behaviors among interacting LLMs (park2023generative; li2023camel) and multi-turn jailbreaking (wei2024jailbroken; chao2023jailbreaking) shows that conversational context reshapes behavior, but focus on creativity or adversarial manipulation rather than failures in well-intentioned (non-adversarial) agents that differ in realistic goals. This paper complements and extends these related works by isolating and quantifying AxA specific behaviors that are invisible in single-agent studies.
@@
This work makes the following contributions: (i) We formalize AxA interactions and introduce echoing, an identity inconsistency failure emergent in such interactions (section 2). (ii) We conduct a large-scale empirical study across providers, domains, and prompts, showing that echoing is prevalent ($5$ to $70\%$), persists in reasoning models ($32.8\%$ on average), and exhibits strong domain sensitivity (section 4.1). (iii) We analyze conversation dynamics, showing that echoing typically arises as interaction grows longer ($7+$ turns in experiments) and that completion-focused metrics do not identity these failures (section 4.3). (iv) We propose a near-term mitigation at the protocol-level that reduces echoing through targeted structured responses (section 1). Our findings highlight echoing is a critical challenge in AxA systems and requires changes to modeling, training, and evaluation approaches tailored to agent-agent interactions.
@@ 2.1 AxA interactions
At turn $t$, agent $A_{i}$ receives its conversational partner’s message $m_{t-1}$ and works through its internal state $s_{t}^{i}$ (its conversation history, thinking, tool outputs, and identity specifications) to construe a response. Specifically, the agent $A_{i}$ uses tool calls ($T_{i}$) and executions to generate an action $a_{t}^{i}$ or generates a response $m_{t}$ that is passed as input to the other agent. Each agent $A_{i}$ is equipped with an end_conversation tool to signal the end of the interaction at any point in an interaction. The LLM policy $\pi_{i}$ operates under a system prompt that encodes the agent’s identity $I_{i}$, objective $O_{i}$, and utility specification $U_{i}$ (see Figure 2). Note that a “turn” in our setup denotes the complete observe–decide–act loop for $A_{i}$ resulting in a response to $m_{t-1}$ and may involve multiple internal LLM calls and tool invocations before committing to an output text. See Appendix B and C for additional details with examples of how this setup is implemented in experiments.
@@ 2.2 AxA Environment
The LLM policy $\pi_{i}$ for agents are implemented through configurable backends spanning OpenAI (GPT-4o, GPT-4.1, o3, GPT-5), Google (Gemini-2.5-Flash, Gemini-2.5-Pro), and Anthropic (Claude Sonnet-4) models, with non-reasoning and reasoning effort controls where applicable. The OpenAI models make use of the responses API (openaiResponses) while the remaining models are accessed via chat completion (openaiChatCompletion).
Each agent’s action space is defined with domain-specific tools $T_{i}$ that include (i) information tools for querying knowledge available (e.g., inventory, pricing) without revealing private constraints, (ii) action tools for persistent environment modification (e.g., bookings, offers), and (iii) communication tools (e.g. end_conversation).
@@ 2.3 AxA Echoing Metric
We presently propose a metric to capture echoing through a domain-specific LLM-based evaluator (zheng2023judging; wang2024direct; gu2024survey) that analyzes the complete conversation history $H_{T}$ for identity inconsistency. The evaluator employs a structured assessment, namely,
@@ 3 Experimental Setup
We evaluate echoing across $60$ AxA configurations spanning $20$ customer-like models, $3$ seller configurations, $3$ domains, and $3$ prompt variants. We focus on customer–seller interactions, treating the customer agent as the primary variable with fewer seller variations to isolate echoing susceptibility. This choice reflects our observation that customer agents are more prone to echoing than seller agents (Figure 3), possibly due to training data distributions that emphasize enterprise roles. This observation is significant as it highlights a fundamental departure from standard LLMs and agents, which when assigned human-facing roles in AxA exhibit new behaviors unseen in other contexts.
@@
Domains. We study three settings with structural goal misalignment. In Hotel booking, customer agents seek optimal accommodations within budget constraints while hotel agents maximize revenue through strategic pricing and upselling. Car sales involves buyer agents evaluating vehicles against budget and requirements (vehicle type, car features) while dealer agents pursue profit maximization. Supply chain procurement features customer agents optimizing cost, quality, and delivery timelines while supplier agents maximize revenue through strategic pricing.
@@
The seller agents is equipped with specific tools to obtain information available to it. For instance: $(i)$ the hotel agent will interact with room inventory databases and pricing systems, $(ii)$ the car dealer will be operating with a vehicle inventory and financing calculation tools, and $(iii)$ the supplier agent will be checking on available product in the inventory.
@@
Each domain further includes domain-specific tools that enable realistic implementation while preserving information asymmetry between agents. A key detail to note is that objectives are often misaligned but the interaction is not strictly zero-sum, i.e, the setups are not adversarial: a higher-priced option (e.g., paying extra for a city-view room) can increase customer agent’s utility which might actually increase the seller agent’s utility as well (higher profit margins). To capture these dynamics, we explicitly separate the objective $O_{i}$ from its utility function $U_{i}$ that determines the value in a completed transaction. These are made transparent to the agent via its system prompt.
@@
Configurations. We test $20$ customer agent models across $7$ models: OpenAI (GPT-4o, GPT-4.1, o3 series, GPT-5 series), Google (Gemini-2.5-Flash, Gemini-2.5-Pro), and Anthropic (Sonnet-4). We set temperature to $0.1$ for non-reasoning model variants and test reasoning models with $3$ effort levels (low, medium, high). The seller agents, on the other hand, were tested with three configurations: GPT-4o (temp $0.1$), GPT-5 (medium reasoning effort), Gemini-2.5-Pro (medium reasoning effort). Each AxA interaction follows turn-based conversation with domain-specific tools until task completion or if each agent has responded atleast $12$ turns. Moreover, each agentic turn is constrained to a maximum of $10$ LLM calls. Results are obtained on at least $10$ independent runs per configuration, yielding approximately $2000$ conversations for validation and analysis.
@@
Prompt Variations. To measure how prompting influences echoing, we report results under three distinct identity prompt variants: minimal (basic role assignment), behavioral (minimal + context + intent), and identity boundary (behavioral + anti-echoing instructions). In the hotel booking domain, this corresponds to a customer agent’s identity $I_{i}$ defined in $3$ levels as (i) You are a customer agent booking a hotel room., (ii) ++ Act as someone who needs accommodation and is evaluating options to find the best value within your constraints., and (iii) ++You represent the customer’s interests only - you are not responsible for the hotel’s business success, revenue, or occupancy goals. Focus exclusively on securing the best booking for your needs. This progression lets us assess whether echoing is a prompt artifact and whether stronger prompts mitigate it. See Appendix C for the complete prompt with identity variations for both agents in a domain.
Evaluation. For echoing detection (equation 1), we make use of GPT-4o with structured response, analyzing complete conversation histories for persona inconsistency as defined in Section 2.3. We note that the analysis is performed only on successfully completed conversations ( representing $93.2\%$ of all generated conversation) to separate behavioral consistency from task performance. To validate the LLM evaluation we performed manual review with stratified sampling (randomly selected equal number of positive and negative cases of echoing) on available conversations. Correlation analysis shows agreement rate of $91.1\%$ with human annotations (see Appendix A).
@@
Figure 3: Echoing rates vs model providers: (Left) Echoing rate is aggregated across all domains and seller agents. Error bars in both plots represent $95\%$ confidence intervals reflecting variance across different model configurations within each category. It is clear that the rate of echoing varies drastically depending on the underlying underlying LLM used for the agent.(Right) Echoing Bias - Percentage of echoing that is attributed to the customer agent vs seller agent per domain aggregated across all agent configs in AxA. We observe that echoing is more prevalent in customer agents.

### v2

@@ Abstract
As large language model (LLM) based agents interact autonomously with one another, a new class of failures emerges that cannot be predicted from single agent performance: behavioral drifts in agent-agent conversations (AxA). Unlike human-agent interactions, where humans ground and steer conversations, AxA lacks such stabilizing signals, making these failures unique. We investigate one such failure, echoing, where agents abandon their assigned roles and instead mirror their conversational partners, undermining their intended objectives. Through experiments across $66$ AxA configurations, $4$ domains (3 transactional, 1 advisory), and $2500+$ conversations (over 250,000250{,}000 LLM inferences), we show that echoing occurs across major LLM providers, with echoing rates as high as $70\%$ depending on the model and domain. Moreover, we find that echoing is persistent even in advanced reasoning models with substantial rates ($32.8\%$) that are not reduced by reasoning efforts. We analyze prompt, conversation dynamics, showing that echoing arises as interaction grows longer ($7+$ agent turns) and is not merely an artifact of sub-optimal experiment design. Finally, we introduce a protocol-level mitigation where targeted use of structured response reduces echoing to $9\%$.
@@ 1 Introduction
Recent advances in large language models (LLMs) Jaech et al., 2024; Comanici et al., 2025; Guo et al., 2025 have enabled agentic systems that can plan, reason, and act in open-ended settings (Mialon et al., 2023; Guo et al., 2024; Wang et al., 2024a). A natural next step is agent-agent (AxA) interaction: agents that converse directly with one another to collaborate, negotiate, and execute tasks on behalf of users or organizations (Raskar et al., 2025; IBM BeeAI, 2025; Google A2A, 2025; Outshift by Cisco, 2025; Tomasev et al., 2025). Although AxA presents a frontier, our understanding of its reliability and requirements is still incomplete; including how consistently agents maintain their roles and objectives across multi-turn interactions.
@@
Existing research and evaluation frameworks are focused on single-agent capabilities, i.e., how well a model performs tasks in isolation or in a human supervised settings (Yao et al., 2024; Huang et al., 2025; Sirdeshmukh et al., 2025). Even when multi-turn dialog is considered, benchmarks typically rely on simulated user (instructions, tools, or context) and measure single-agent success criteria (Barres et al., 2025). These task driven setups are unable to capture behaviors that emerge specifically from AxA, where agents with private, potentially misaligned, objectives must interact, namely, an user agent that is not simulated only to provide information for task completion. In contrast to human-agent interactions, where human feedback is accumulated continually, and often subtly, steering and grounding behavior, AxA relies on predefined instructions/specifications and an often untested assumption of alignment between agents (see Figure 1).
@@
This paper studies an AxA specific failure mode that we call echoing: an agent abandons its assigned identity and mirrors its conversational partner, undermining its stated objectives. We provide a systematic investigation of this behavioral failure across several models, domains, and prompts. Through experiments spanning $66$ configurations, $4$ domains ($3$ transactional: car sales, hotel booking, supply chain; $1$ advisory: medical consultation), $2500+$ conversations, we find that echoing occurs across major providers with rates as high as $70\%$ depending on the model and domain. Echoing persists even in advanced reasoning models (average $32.8\%$) and does not diminish with increased reasoning effort; non-reasoning variants average $37.7\%$ (section 4.2.1). Our empirical results shows that prompt engineering reduces but does not eliminate the failure (section 4.2.2), suggesting a limitation akin to hallucination but specific to AxA. We further observe that task completion metrics mask these failures as $93\%$ of conversations completed successfully, even when identity drift occurred. Further, the outcome value varied substantially within a single setup.
@@
Figure 1: Agent x Agent setup: (Left) Two agents, a customer agent and a seller agent, given instructions, objectives, private tools, and resources, are entrusted to complete a particular task given a situation-specific spec. The customer agent, in this case, negotiates a room on behalf of a human (with specific requirements) with a seller agent, a hotel agent representing an enterprise with specifics pertinent to the hotel. (Right) Conversation snippet from an AxA exchange where the customer agent echoes the language and behavior more appropriate of an hotel agent. The seller agent, in this example, continued the interaction without correction and ended up accepting the package proposed by the customer agent. Such a failure is unlikely in human-agent interactions and even when it arises would typically be corrected by the human ensuring that the agent remains aligned with its intended role. More examples of echoing are provided in Appendix D.
@@
Our setting differs from multi-agent systems (MAS) (Stone and Veloso, 2000; Tampuu et al., 2017) in that the AxA agents maintain private state, operate with distinct tools, and may have competing utilities. In contrast, MAS emphasize coordinated agent teams with shared goals, centralized orchestration, and even shared states; recent LLM-based MAS continues this paradigm for task decomposition and parallelism (Wu et al., 2023; Hong et al., 2023; Chen et al., 2024). AxA focus is autonomous conversational entities. In conversational AI, existing works study human-agent interaction where humans provide feedback and course correction (Serban et al., 2017; Roller et al., 2020). Moreover, post-training, alignment techniques are optimized for human-facing use (Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2023; Song et al., 2024) and not AxA (might even be cause of biases such as over-accommodation, role drifts). Work on alignment, both objective adherence (Christiano et al., 2017; Russell, 2019) and multi-agent coordination (Carroll et al., 2019; Eccles et al., 2019), rarely addresses an AxA setting. Other works on emergent behaviors among interacting LLMs (Park et al., 2023; Li et al., 2023) and multi-turn jailbreaking (Wei et al., 2024; Chao et al., 2023) shows that conversational context reshapes behavior, but focus on creativity or adversarial manipulation rather than failures in well-intentioned (non-adversarial) agents that differ in realistic goals. This paper complements and extends these works by isolating and quantifying AxA specific behaviors that are invisible in single-agent studies (see Appendix H for more details).
@@
This work makes the following contributions: (i) We formalize AxA interactions and introduce echoing, an identity inconsistency failure emergent in such interactions (section 2). (ii) We conduct a large-scale empirical study across providers, domains, and prompts, showing that echoing is prevalent ($5$ to $70\%$), persists in reasoning models ($32.8\%$ on average), and exhibits strong domain sensitivity (section 4.1). (iii) We analyze conversation dynamics, showing that echoing typically arises as interaction grows longer ($7+$ turns in experiments) and that completion-focused metrics do not identity these failures (section 4.3). (iv) We explore first-reflex mitigation strategies (prompting, reasoning, protocol-level interventions), demonstrating that while these attenuate echoing, they cannot eliminate it—consistent with the hypothesis that echoing stems from fundamental mismatch in models rather than suboptimal deployment choices (section 4.2, section 4.4). Our findings highlight echoing is a critical challenge in AxA systems and requires changes to modeling, training, and evaluation approaches tailored to agent-agent interactions.
@@ 2.1 AxA interactions
At turn $t$, agent $A_{i}$ receives its conversational partner’s message $m_{t-1}$ and works through its internal state $s_{t}^{i}$ (its conversation history, thinking, tool outputs, and identity specifications) to construe a response. Specifically, the agent $A_{i}$ uses tool calls ($T_{i}$) and executions to generate an action $a_{t}^{i}$ or generates a response $m_{t}$ that is passed as input to the other agent. Each agent $A_{i}$ is equipped with an end_conversation tool to signal the end of the interaction at any point in an interaction. The LLM policy $\pi_{i}$ operates under a system prompt that encodes the agent’s identity $I_{i}$, objective $O_{i}$, and utility specification $U_{i}$ (see Figure 2). Note that a “turn” in our setup denotes the complete observe-decide-act loop for $A_{i}$ resulting in a response to $m_{t-1}$ and may involve multiple internal LLM calls and tool invocations before committing to an output text. See Appendix B and C for additional details with examples of how this setup is implemented in experiments.
@@ 2.2 AxA Environment
The LLM policy $\pi_{i}$ for agents are implemented through configurable backends spanning OpenAI (GPT-4o, GPT-4.1, o3, GPT-5), Google (Gemini-2.5-Flash, Gemini-2.5-Pro), Anthropic (Claude Sonnet-4), and Meta (Llama-3.1-8B, Llama-3.1-70B) models, with non-reasoning and reasoning effort controls where applicable. The OpenAI models make use of the responses API (OpenAI, 2025) while the remaining models are accessed via chat completion (OpenAI, 2023).
Each agent’s action space is defined with domain-specific tools $T_{i}$: (i) information tools for querying knowledge (e.g., inventory) without revealing private constraints, (ii) action tools for persistent environment modification (e.g., bookings), and (iii) communication tools (e.g. end_conversation).
@@ 2.3 AxA Echoing Metric
We presently propose a metric to capture echoing through a domain-specific LLM-based evaluator (Zheng et al., 2023; Wang et al., 2024b; Gu et al., 2024) that analyzes the complete conversation history $H_{T}$ for identity inconsistency. The evaluator employs a structured assessment, namely,
@@ 3 Experimental Setup
We evaluate echoing across $66$ AxA configurations spanning $22$ customer-like models, $3$ seller configurations, $3$ transactional domains, and $3$ prompt variants. We focus on transactional customer-seller interactions as these represent immediate deployment contexts for AxA systems, with multiple industry initiatives targeting such scenarios. Moreover, transactional domains provide well-defined role boundaries that make echoing behavior unambiguous to detect and measure. Results with a non-transactional domain (medical consultation) are provided for completeness. We treat the customer agent as the primary variable with fewer seller variations to isolate echoing susceptibility. This choice reflects our observation that customer agents are more prone to echoing than seller agents (Figure 3), possibly due to training data distributions that emphasize enterprise roles. When assigned consumer-facing roles in AxA, these models exhibit behavior unseen in human-facing training contexts. Additional rationale for our experimental design choices, including domain selection and the asymmetric customer-seller configuration, is provided in Appendix G.
@@
Transactional Domains. We study three transactional settings with structural goal misalignment. In Hotel booking, customer agents seek optimal accommodations within budget constraints while hotel agents maximize revenue through strategic pricing and upselling. Car sales involves buyer agents evaluating vehicles against budget and requirements (vehicle type, car features) while dealer agents pursue profit maximization. Supply chain procurement features customer agents optimizing cost, quality, and delivery timelines while supplier agents maximize revenue through strategic pricing.
@@
The seller agents are equipped with tools to obtain information available to it, i.e, $(i)$ the hotel agent will interact with room inventory and pricing , $(ii)$ the car dealer will operate with a vehicle inventory and financing, and $(iii)$ the supplier agent will check on available product in inventory.
@@
Each domain is set to enable realistic implementation while preserving information asymmetry between agents. Note that objectives in AxA are often misaligned but the interaction is not strictly zero-sum, i.e, the setups are not adversarial: a higher-priced option (e.g., paying extra for a city-view room) can increase customer agent’s utility which might actually increase the seller agent’s utility as well (higher profit margins). To capture these dynamics, we explicitly separate the objective $O_{i}$ from its utility function $U_{i}$ that determines the value in a completed transaction. These are made transparent to the agent via its system prompt.
@@
Configurations. We test $22$ customer agent models across OpenAI (GPT-4o, GPT-4.1, o3 series, GPT-5 series), Google (Gemini-2.5-Flash, Gemini-2.5-Pro), Anthropic (Sonnet-4), and Meta (Llama-3.1-8B, Llama-3.1-70B). We set temperature to $0.1$ for non-reasoning model variants and test reasoning models with $3$ effort levels (low, medium, high). The seller agents, on the other hand, were tested with three configurations: GPT-4o (temp $0.1$), GPT-5 (medium reasoning), Gemini-2.5-Pro (medium reasoning). Each AxA interaction follows turn-based conversation with domain-specific tools until task completion or if each agent has responded atleast $12$ turns. Each agentic turn is constrained to a maximum of $10$ LLM calls. Results are obtained on at least $10$ independent runs per configuration, yielding over $2500$ conversations for validation and analysis. In total over 250,000250{,}000 LLM inferences were made for the analysis excluding judge evaluations and development iterations.
@@
Prompt Variations. To measure how prompting influences echoing, we report results under three distinct identity prompt variants: minimal (basic role assignment), behavioral (minimal + context + intent), and identity boundary (behavioral + anti-echoing instructions). In the hotel booking domain, this corresponds to a customer agent’s identity $I_{i}$ defined in $3$ levels as (i) You are a customer agent booking a hotel room., (ii) ++Act as someone who needs accommodation and is evaluating options to find the best value within your constraints., and (iii) ++You represent the customer’s interests only - you are not responsible for the hotel’s business success, revenue, or occupancy goals. Focus exclusively on securing the best booking for your needs. This progression lets us assess whether echoing is a prompt artifact and whether stronger prompts mitigate it. See Appendix C for details.
@@
Figure 3: Echoing rates vs model providers: (Left) Echoing rate is aggregated across transactional domains, seller agents, and all configurations (both reasoning and non-reasoning model variants). Error bars in both plots represent $95\%$ confidence intervals reflecting variance across different model configurations within each category. It is clear that the rate of echoing varies drastically depending on the underlying LLM used for the agent.(Right) Echoing Bias - Percentage of echoing that is attributed to the customer agent vs seller agent per domain aggregated across all agent configs in AxA. We observe that echoing is more prevalent in customer agents.
Evaluation. For echoing detection (equation 1), we make use of GPT-4o with structured response, analyzing complete conversation histories for persona inconsistency as defined in Section 2.3. We note that the analysis is performed only on successfully completed conversations (representing $93.2\%$ of all AxA runs) to separate behavioral consistency from task performance. A conversation is considered “successful” if a transaction is completed (e.g., booking made, car sold, contract agreed). Critically, echoing does not prevent task completion but undermines the quality and representativeness of outcomes. For instance, a customer agent echoing the seller results in a completed hotel booking but at sub-optimal terms for one of the parties involved. This demonstrates how standard success metrics mask behavioral failures. To validate the LLM evaluation we performed manual review on $150$ conversation obtained with stratified sampling (randomly selected equal number of positive and negative cases of echoing for each domain). Correlation analysis shows agreement rate of $91.1\%$ with human annotations (see Appendix A for ablations and additional details).